OpenStudy (anonymous):

**Advance Algebra With Financial Application** [WORK SHOWN] Arnold is financing a condominium for $592,000. He gets a 30-year, fixed rate mortgage with a rate of 5.575%. How much interest will he pay over the life of his loan? Round to the nearest dollar. $1,220,119 $43,585 $181,614 $628,119 ---------- * MY WORK * M=592000 [(0.0046458333)(1+0.0046458333)^(12*30)/(1+0.0046458333)^(12*30)-1]

OpenStudy (anonymous):

Well we need to find the value of the monthly payments multiply. Then with the monthly payments we need to multiply that by 360 months because you pay monthly for 30 years. Once you find that just subtract the principle and you will find the total amount of interest you will pay over the life of the loan.

OpenStudy (anonymous):

@mebs -589.250 ?

OpenStudy (anonymous):

hmm you did something wrong lol it can't be a negative number....or can it !!!!

OpenStudy (anonymous):

2749.33 ?

OpenStudy (anonymous):

was the formula \[M =\frac{ P(1+(1-i))^{-n} }{ i}\]

OpenStudy (anonymous):

I have trouble recalling it do you have it... if you write it down please.

OpenStudy (anonymous):

just plug in your variables correctly principle = 592 000 interest rate = 0.0575 n = 30 compounding period 12 so i = 0.0575/12 = 0.004791666 you n = 12 * 30 = 360

OpenStudy (anonymous):

remember the exponent part is to the -n

OpenStudy (anonymous):

M=592000(1+(1-0.004791666)/(0.004791666) ?

OpenStudy (anonymous):

no my formula is wrong I am looking for the correct one...

OpenStudy (anonymous):

Here it is I think this is the right formula. \[PV= \frac{ R[ 1-(1+i)^{-n}] }{ i }\] Where PV is the present value of the condo and R is the monthly payments.

OpenStudy (anonymous):

Let me use the one I used I grade 11.... So the interest rate is 0.05575 n = 30 years comound (c) =12 so interest = 0.004645833 the periods = 12 * 30 = 360 The formula for the Monthly payments M. \[M = \frac{ P \times i }{ [(c)(1-(1+\frac{ i }{ c })^{-nc}] }\] Substitute the values. \[M = \frac{ 592000 \times 0.005575 }{ [(12)(1-(1+\frac{ 0.005575 }{ 12 })^{-(30\times12)}] }\] Your monthly payment should turn out to be $2231.8797 rounded to $2231.88 now multiply it by 360. ( I used the non rounded number $803, 476.70. If you use the rounded number you get $803,476.80. Now to find the interest just subtract $592 000. So the amount of interest you pay is: $211 476.80 . Sorry for the confusion it has been a while since I took that class.